# Revisiting Reweighted Risk for Calibration: AURC, Focal Loss, and Inverse Focal Loss

### v2

@@ 4 A Differentiable Loss Function Derived from Selective Classification
In this section, we begin by examining lower bounds for two widely used calibration metrics—ECE and class-wise ECE—as presented in Prop. 4.14.3. These results provide theoretical insight into how the calibration error can be reduced for a pre-trained model and motivate a deeper exploration of the connection between selective classification and calibration. Building on these insights, we propose a differentiable loss function derived from the AURC metric in selective classification. The formulation is fully differentiable—enabled by the SoftRank method—and supports arbitrary CSFs.
@@ 4.1 Linking Calibration Error Minimization to Selective Classification
Lemma 4.1 (Lower bound of ECE [36, Prop. 2]).
Let $f:\mathcal{X}\!\to\!\Delta^{k}$ be the classifier and $\widehat{\operatorname{err}}(f)$ be its empirical error rate on a finite set $\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\}_{i=1}^{n}$ given by $\widehat{\operatorname{err}}(f)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\bigl{(}\hat{y}_{i}\neq y_{i}^{\prime}\bigr{)}$ where $\hat{y}_{i}=\arg\max\limits_{c\in[k]}f_{c}(x_{i})$. For the top-label ECE computed over all binning schemes $B\in\mathcal{B}$ holds that
@@
In the proof of Prop. 4.1, Ma and Blaschko [36] demonstrate that a calibrated model can be constructed using a binary classifier that separates correct from incorrect predictions. We reinterpret this insight through the lens of selective classification, yielding the following result.
Proposition 4.2 (Optimal Calibration Map for ECE [36, Prop. 3]).
@@
$\sup_{B\in\mathcal{B}}\widehat{\operatorname{ECE}}_{B}(f^{\prime})\leq\sup_{B\in\mathcal{B}}\widehat{\operatorname{ECE}}_{B}(f),$ (14)
@@
$\max_{c}f^{\prime}_{c}(\boldsymbol{x})=\begin{cases}1&\text{if }g(\boldsymbol{x})\geq\tau,\\ \frac{1}{k}&\text{if }g(\boldsymbol{x})<\tau.\end{cases}$ (15)
@@
This proposition suggests that ECE can be improved by boosting the predicted confidence for correctly classified samples while reducing it for incorrect ones. Notably, assigning a uniform prediction $\frac{1}{k}$ when $g(\boldsymbol{x})<\tau$ mimics the abstention mechanism used in selective classification. In this context, the CSF with a threshold $\tau$ effectively acts as a binary classifier. To further support the connection between calibration errors and selective classification, we next extend Prop. 4.1 to the cwECE and derive a constructive rule for minimizing it.
@@
$\operatorname{cwECE}(f)\geq\sup_{B\in\mathcal{B}}\widehat{\operatorname{cwECE}}_{B}(f)>\frac{1}{k}\widehat{\operatorname{err}}(f).$ (16)
The proof is provided in App. 7.2, and we sketch it here. First, we apply Minkowski’s inequality to derive a lower bound on the supremum over all class-wise binning schemes. Next, we show that this supremum is bounded below by $\widehat{\operatorname{err}}(f)/k$, unless $f$ is perfectly accurate. Finally, by invoking Jensen’s inequality, as in Kumar et al. [30], the lower bound follows. This construction naturally leads to a procedure for improving cwECE, analogous to Prop. 4.2.
@@
$\sup_{B\in\mathcal{B}}\widehat{\operatorname{cwECE}}_{B}(f^{\prime})\leq\sup_{B\in\mathcal{B}}\widehat{\operatorname{cwECE}}_{B}(f)$ (17)
with
$\max_{c}f^{\prime}_{c}(\boldsymbol{x})=\begin{cases}1&\text{if }g(\boldsymbol{x})\geq\tau,\\ \min\left(\frac{1}{2},\max_{c}f_{c}(\boldsymbol{x})\right)&\text{otherwise}.\end{cases}$ (18)
@@
This construction rule follows directly from the proof of Prop. 4.3. While slightly different from Prop. 4.2, this result similarly suggests that reducing the maximum predicted probability for uncertain samples helps improve cwECE. Together, these propositions build a theoretical foundation connecting selective classification with calibration error minimization: while Prop. 4.2 addresses ECE, Prop. 4.3 and its constructive variant extend this idea to cwECE. We conclude by presenting a result from Gruber and Buettner [16], which synthesizes calibration metrics under a unified theoretical bound.
Proposition 4.5 (Hierarchy Among Calibration Metrics [16, Theorem 3.1]).
Let $f:\mathcal{X}\to\Delta^{k}$ be a classifier, and let $1\leq\rho\leq 2$. Then,
$n^{\frac{1}{\rho}-\frac{1}{2}}\sqrt{\text{BS}(f)}\geq\text{CE}_{\rho}(f)\geq\begin{cases}k\cdot\operatorname{cwECE}_{\rho}(f),\\ \operatorname{ECE}(f),\end{cases}\geq 0,$ (19)
@@
4.2 Regularized AURC Loss Function
Given the preceding results, reducing calibration error requires lowering the maximum predicted probability for low-confidence samples—a strategy that echoes the "abstain" mechanism central to selective classification. Since the AURC metric originates from the selective classification framework and employs a weighting scheme similar to that of inverse focal loss (see Fig. 1), such reweighting schemes are theoretically well-suited for calibration error minimization. Motivated by this, we propose the regularized AURC (r-AURC) loss to examine whether this reweighting approach can effectively enhance model calibration. An asymptotically unbiased Monte Carlo estimator [47] for the population AURC defined in Eq. (7) is given by
$\widehat{\operatorname{AURC}}_{g}(f)=-\frac{1}{n}\sum_{i=1}^{n}\ln\Big{(}1-\frac{r_{i}}{n+1}\Big{)}\ell(\boldsymbol{p}_{i},\boldsymbol{y}_{i}),$ (20)
where $r_{i}:=r(g(\boldsymbol{p}_{i}))$ denotes the rank of the confidence score $g(\boldsymbol{p}_{i})$ among all samples in ascending order. Thus, a larger $r_{i}$ corresponds to a higher confidence. This estimator is consistent and can be directly used as a training loss to minimize the AURC metric. Importantly, since $r_{i}=r(g(\boldsymbol{p}_{i}))$ is a function of the model output $\boldsymbol{p}_{i}$, its gradient $\frac{\partial r_{i}}{\partial\boldsymbol{p}_{i}}$ must be considered in backpropagation. To facilitate end-to-end optimization, we adopt the SoftRank approximation [3], which replaces the hard (non-differentiable) rank with a smooth, differentiable surrogate $\hat{r}_{i}$. This enables standard gradient-based training, as outlined in Algorithm 1.
Algorithm 1 AURC Optimization with $\mathrm{SoftRank}$
@@ 4.2 Regularized AURC Loss Function
1:Training batch $\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\}_{i=1}^{n}$, model $f_{\theta}$, a CSF $g(\cdot)$, learning rate $\eta$
2:Update: $\theta\leftarrow\theta-\eta\nabla_{\theta}L$
3:Forward Pass:
4:$\boldsymbol{p}_{i}\leftarrow f_{\theta}(\boldsymbol{x}_{i})$triangleright Model prediction
5:$\hat{r}_{i}\leftarrow\mathrm{SoftRank}(g(\boldsymbol{p}_{i}))$
6:$L\leftarrow\frac{1}{n}\sum_{i=1}^{n}\left[-\ln(1-\hat{r}_{i})\cdot\ell(\boldsymbol{p}_{i},\boldsymbol{y}_{i})\right]$triangleright AURC loss
@@
9:     Compute $\nabla_{\boldsymbol{p}_{i}}L$ via Eq. (41)
10:     $\nabla_{\theta}L\mathrel{+}=(\nabla_{\boldsymbol{p}_{i}}L)^{\top}\nabla_{\theta}\boldsymbol{p}_{i}$
@@
This loss function typically relies on sorting based on the confidence score function (CSF), incurring a computational cost of $\mathcal{O}(n\log(n))$, which is generally more efficient than kernel-based calibration error losses that require quadratic time. To better understand the effect of CSF on AURC metric, we define the perfectly aligned CSF.
Definition 4.6 (Perfectly Aligned CSF).
A CSF $g^{\star}$ is said to be perfectly aligned with the loss function $\ell$ if, for any two predictions $\boldsymbol{p}_{i}$, $\boldsymbol{p}_{j}$ with corresponding labels $\boldsymbol{y}_{i}$, $\boldsymbol{y}_{j}$, the following holds:
$g^{\star}(\boldsymbol{p}_{i})\leq g^{\star}(\boldsymbol{p}_{j})\quad\Rightarrow\quad\ell(\boldsymbol{p}_{i},\boldsymbol{y}_{i})\geq\ell(\boldsymbol{p}_{j},\boldsymbol{y}_{j}).$ (21)
This condition ensures that higher confidence scores correspond to lower losses, meaning that the confidence ranking is consistent with the loss. In practice, this alignment significantly affects the effectiveness of the r-AURC loss, as it governs how the sample weights (via soft rank) interact with the loss values.
Proposition 4.7 (Lower Bound for Empirical AURC).
Let $g^{\star}$ be a CSF that is perfectly aligned with the loss function $\ell$. Let $\mathcal{G}$ denote the set of all valid confidence mappings. For a given model $f$ and the cross-entropy loss, the empirical AURC is minimized at $g^{\star}$, and has the following lower bound:
$\inf_{g\in\mathcal{G}}\widehat{\text{AURC}}_{g}(f)=\widehat{\text{AURC}}_{g^{\star}}(f)\geq\frac{n\log 2}{2(n+1)}\widehat{\operatorname{err}}^{2}(f)+\frac{\log 2}{2(n+1)}\widehat{\operatorname{err}}(f),$ (22)
where $\widehat{\operatorname{err}}(f)$ is the empirical error of $f$ on a finite dataset.
See App. 7.2 for the proof. This proposition indicates that AURC is minimized when the CSFs are perfectly aligned with the loss and when the model achieves high predictive accuracy. In addition, Prop. 4.1 and 4.3 show that lower calibration error also relies on the model being accurate. These insights collectively suggest that both calibration and accuracy are mutually reinforcing goals.
To this end, we consider the following constrained optimization framework to learn a model $f$ that is both accurate and well-calibrated: $f^{*}=\arg\min_{f\in\mathcal{F}}\mathcal{R}(f),\quad\text{subject to}\quad\text{AURC}_{g}(f)\leq B,$ where $\mathcal{R}(f)$ denotes the expected predictive risk, and $B>0$ is a user-defined budget. The corresponding Lagrangian relaxation of this constrained objective is given by:
$f^{*}=\arg\min_{f\in\mathcal{F}}\left((1-\lambda)\,\mathcal{R}(f)+\lambda\,\text{AURC}_{g}(f)\right),$ (23)
where $\lambda\in[0,1]$ controls the trade-off between accuracy and AURC minimization. Note that the behavior of this regularized loss depends on the choice of the CSF $g$ used in the AURC. As we show in later sections, different CSFs may lead to different calibration-performance.
@@ 5 Experiments
Datasets and Models. We conduct experiments on image classification datasets including CIFAR-10, CIFAR-100 [27], and Tiny-ImageNet [8]. For Tiny-ImageNet, all images are resized to 224×224 pixels. On CIFAR-10/100, we evaluate models including WideResNet-28x10 [46], PreResNet-56 [21], and ResNet architectures with depths 50 and 110 [20]. For Tiny-ImageNet, we use pre-trained models from the timm library [45]. Each experiment is repeated with five different random seeds, and we report the mean of the resulting performance metrics.
@@
Evaluation Metrics. We evaluate calibration performance using the binned estimators of ECE and class-wise ECE (cwECE), computed under two binning strategies: equal-width binning (denoted as ECE and cwECE) and adaptive binning [9] (denoted as ECEa and cwECEa). In addition, we consider a kernel density estimation-based calibration error, denoted as $\mathrm{ECE}^{\mathrm{KDE}}$ [41], which provides a continuous, non-binned alternative. To assess whether our proposed r-AURC loss directly improves its intended objective, we also report the AURC metric in all evaluations.
Baselines. We explore the trainable calibration methods including cross-entropy loss (XE), KDE-XE [41], focal loss [37], inverse focal loss [43], and dual focal loss [42]. For focal loss, we consider FL-53 where $\gamma=5$ for $f_{y^{\prime}}(\boldsymbol{x})\in[0,0.2)$ and $\gamma=3$ otherwise. For the latter two loss functions, the $\gamma$ parameter is fine-tuned on the split validation set following the settings in Tao et al. [42]. For inverse focal loss, we note that when $\gamma>3.5$, it may not able to update the model properly due to the positive gradients, as detailed discussed in App. 7.3. For KDE-XE, we set its regularization parameter $\lambda=0.3$ on CIFAR-10, and $\lambda=0.5$ on CIFAR-100, to achieve the best canonical calibration error as suggested in the original paper. We also evaluate post-hoc calibration methods, such as MetaAcc [36] and Temperature Scaling (TS) [17], on models trained with different loss functions. (See App. 7.5 for additional results on those baselines.)
Table 1: Test accuracy (%) and classwise ECEa ($\downarrow$), reported as mean ×$\times 10^{2}$ over 5 seeds.
Accuracy (%) $\uparrow$ cwECEa ($\downarrow$)
Model XE FL-53 Inv.FL r-Ap r-Ae r-Am XE FL-53 Inv.FL r-Ap r-Ae r-Am
CIFAR-10
ResNet50 94.75 94.43 95.00 94.61 94.30 94.63 0.266 0.591 0.287 0.260 0.249 0.259
ResNet110 94.56 94.32 95.05 94.44 94.25 94.53 0.275 0.544 0.294 0.264 0.271 0.269
WRN 95.30 94.03 95.42 95.42 95.17 95.56 0.236 1.059 0.256 0.251 0.244 0.220
PreResNet56 93.39 93.05 93.78 93.36 93.22 93.38 0.308 0.592 0.383 0.324 0.323 0.313
CIFAR-100
ResNet50 77.54 68.08 77.92 77.68 76.30 77.62 0.092 0.239 0.100 0.091 0.091 0.091
ResNet110 77.21 71.13 77.32 77.43 71.87 77.49 0.098 0.178 0.111 0.097 0.131 0.095
WRN 78.68 72.68 78.73 78.73 78.66 78.86 0.094 0.190 0.106 0.092 0.094 0.098
PreResNet56 72.21 66.06 72.63 72.43 72.53 72.34 0.107 0.192 0.113 0.105 0.090 0.106
Tiny-ImageNet
ResNet50 81.84 78.24 79.56 82.11 82.47 82.31 0.126 0.177 0.136 0.127 0.122 0.123
Results. Table 1 presents the test performance in terms of cwECE across different loss functions. Overall, the proposed r-AURC loss consistently outperforms other baselines, while focal loss tends to yield the worst cwECE in most cases. Notably, the performance of r-AURC is sensitive to the choice of CSF; for instance, MSP proves suboptimal on CIFAR-10, underscoring the importance of selecting a suitable CSF to ensure reliable calibration. In Fig. 2, focal loss tends to be overconfident on easy samples, whereas AURC-based losses exhibit better-calibrated behavior. As shown in Table 2, the cwECEa of r-AURC can be further reduced via TS, highlighting the benefit of post-hoc calibration. Additional results in Appendix 7.5 show that MetaAcc can further enhance cwECE—particularly for losses other than focal loss—though this often comes at the cost of reduced predictive accuracy. Regarding ECE, r-AURC achieves the best overall performance on CIFAR-10 across all architectures, and its ECE can likewise be improved through TS. Notably, focal loss performs relatively well in terms of ECE on CIFAR-100, suggesting its utility in top-label settings. Given that focal loss lacks a direct theoretical connection to calibration error minimization, we advocate for the use of inverse focal loss or r-AURC, both of which are explicitly designed to optimize calibration. Our r-AURC loss does not consistently achieve the lowest ECE across all datasets, suggesting a degree of dataset dependency or the need for incorporating a more reliable CSF to fully realize its calibration potential.
Table 2: Mean cwECE ×$\times 10^{2}$ ($\downarrow$) before and after TS. The temperature $T$ selected by minimizing the ECE on the training set is shown in brackets next to each TS result.
Model XE FL-53 Inv. FL r-Ap r-Ae r-Am
CIFAR-10
ResNet50 0.641 0.838 0.805 0.629 0.621 0.644
     +TS 0.456 [1.30] 0.911 [1.02] 0.692 [1.36] 0.438 [1.29] 0.407 [1.28] 0.457 [1.30]
ResNet110 0.670 0.767 0.884 0.691 0.656 0.671
     +TS 0.478 [1.30] 0.870 [1.03] 0.803 [1.38] 0.478 [1.31] 0.424 [1.29] 0.463 [1.30]
WRN 0.623 1.215 0.701 0.605 0.608 0.588
     +TS 0.417 [1.33] 1.093 [0.97] 0.554 [1.37] 0.407 [1.33] 0.393 [1.33] 0.381 [1.34]
PreResNet56 0.639 0.799 0.881 0.606 0.586 0.628
     +TS 0.462 [1.20] 0.634 [0.95] 0.729 [1.30] 0.424 [1.20] 0.428 [1.18] 0.444 [1.21]
CIFAR-100
ResNet50 0.248 0.285 0.268 0.245 0.246 0.246
     +TS 0.189 [1.36] 0.284 [1.02] 0.200 [1.37] 0.192 [1.36] 0.191 [1.30] 0.189 [1.36]
ResNet110 0.259 0.244 0.293 0.257 0.276 0.252
     +TS 0.194 [1.36] 0.255 [1.07] 0.209 [1.38] 0.191 [1.36] 0.224 [1.27] 0.188 [1.36]
WRN 0.235 0.254 0.267 0.238 0.218 0.240
     +TS 0.189 [1.37] 0.285 [1.09] 0.193 [1.38] 0.191 [1.36] 0.204 [1.35] 0.190 [1.36]
PreResNet56 0.271 0.271 0.308 0.264 0.223 0.265
     +TS 0.228 [1.13] 0.273 [0.96] 0.255 [1.18] 0.220 [1.14] 0.196 [1.10] 0.221 [1.14]
Tiny-ImageNet
ResNet50 0.127 0.177 0.136 0.127 0.122 0.123
     +TS 0.152 [1.21] 0.195 [1.07] 0.160 [1.22] 0.153 [1.21] 0.139 [1.17] 0.151 [1.21]
0.00.20.40.60.81.00.00.20.40.60.81.0ConfidenceAccuracyPerfect CalibrationOutputGapECE=2.39
(a) r-AURCm
0.00.20.40.60.81.00.00.20.40.60.81.0ConfidenceAccuracyPerfect CalibrationOutputGapECE=4.02

### v3

@@ 4 A Differentiable Loss Function Derived from Selective Classification
In this section, we begin by examining lower bounds for two widely used calibration metrics—ECE and cwECE—as presented in Prop. 4.14.3. These results provide theoretical insight into how calibration error can be reduced for a pre-trained model and motivate a deeper exploration of the connection between selective classification and calibration. Building on these insights, we propose a differentiable loss function derived from selective classification. The formulation is fully differentiable—enabled by a bin-based CDF approximation and supports arbitrary CSFs.
@@ 4.1 Linking Calibration Error Minimization to Selective Classification
Lemma 4.1 (Lower bound of ECE (Ma and Blaschko,, 2021, Prop. 2)).
Let $f:\mathcal{X}\!\to\!\Delta^{k}$ be the classifier and $\widehat{\operatorname{err}}(f)$ be its empirical error rate on a finite set $\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\}_{i=1}^{n}$ given by $\widehat{\operatorname{err}}(f)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\bigl(\hat{y}_{i}\neq y_{i}^{\prime}\bigr)$ where $\hat{y}_{i}=\arg\max\limits_{c\in[k]}f_{c}(x_{i})$. For the top‑label ECE computed over all binning schemes $B\in\mathcal{B}$ holds that
@@
In the proof of Prop. 4.1, Ma and Blaschko, (2021) demonstrate that a calibrated model can be constructed using a binary classifier that separates correct from incorrect predictions. We reinterpret this insight through the lens of selective classification, yielding the following result.
Proposition 4.2 (Optimal Calibration Map for ECE (Ma and Blaschko,, 2021, Prop. 3)).
@@
$\sup_{B\in\mathcal{B}}\widehat{\operatorname{ECE}}_{B}(f^{\prime})\leq\sup_{B\in\mathcal{B}}\widehat{\operatorname{ECE}}_{B}(f),$
@@
$\max_{c}f^{\prime}_{c}(\boldsymbol{x})=\begin{cases}1&\text{if }g(\boldsymbol{x})\geq\tau,\\ \frac{1}{k}&\text{if }g(\boldsymbol{x})<\tau.\end{cases}$
@@
This proposition suggests that ECE can be improved by boosting the predicted confidence for correctly classified samples while reducing it for incorrect ones. Notably, assigning a uniform prediction $\frac{1}{k}$ when $g(\boldsymbol{x})<\tau$ mimics the abstention mechanism in selective classification. In this context, the CSF with a threshold $\tau$ effectively acts as a binary classifier. To further support the connection between calibration errors and selective classification, we next extend Prop. 4.1 to the cwECE and derive a constructive rule for minimizing it.
@@
$\operatorname{cwECE}(f)\geq\sup_{B\in\mathcal{B}}\widehat{\operatorname{cwECE}}_{B}(f)>\frac{1}{k}\widehat{\operatorname{err}}(f).$
Proof.
See Appendix B.1. ∎
This construction naturally leads to a procedure for improving cwECE, analogous to Prop. 4.2.
@@
$\sup_{B\in\mathcal{B}}\widehat{\operatorname{cwECE}}_{B}(f^{\prime})\leq\sup_{B\in\mathcal{B}}\widehat{\operatorname{cwECE}}_{B}(f)$
$\text{with }\max_{c}f^{\prime}_{c}(\boldsymbol{x})=\begin{cases}1&\text{if }g(\boldsymbol{x})\geq\tau,\\ \min\left(\frac{1}{2},\max_{c}f_{c}(\boldsymbol{x})\right)&\text{otherwise}.\end{cases}$
@@
This construction rule follows directly from the proof of Prop. 4.3. While slightly different from Prop. 4.2, this result similarly suggests that reducing the maximum predicted probability for uncertain samples helps improve cwECE. Together, these propositions build a theoretical foundation connecting selective classification with calibration error minimization: while Prop. 4.2 addresses ECE, Prop. 4.3 and its constructive variant extend this idea to cwECE. We conclude by presenting a result from Gruber and Buettner, (2022), which synthesizes calibration metrics under a unified theoretical bound.
Proposition 4.5 (Hierarchy Among Calibration Metrics (Gruber and Buettner,, 2022, Theorem 3.1)).
Let $f:\mathcal{X}\to\Delta^{k}$ be a classifier, and let $1\leq\rho\leq 2$. Then:
$n^{\frac{1}{\rho}-\frac{1}{2}}\sqrt{\text{BS}(f)}\geq\text{CE}_{\rho}(f)\geq\begin{cases}k\cdot\operatorname{cwECE}_{\rho}(f),\\ \operatorname{ECE}(f),\end{cases}\geq 0$
@@
4.2 Selective Risk Loss
Given the preceding results from Prop. 4.2 and Prop. 4.4, reducing calibration error requires lowering the maximum predicted probability for incorrectly predicted samples and increasing it for correctly predicted ones. The latter is typically handled by the loss function during training. However, the former has received less attention in trainable calibration methods, and this is the issue we aim to address in this paper from the perspective of selective classification. Motivated by this observation, we propose a loss function similar to AURC, but excludes selective risk over the high-confidence region as these samples are well-calibrated by the loss function during training. Let us denote the low confidence region $\mathcal{X}_{\tau}$ as $\mathcal{X}_{\tau}=\{\boldsymbol{x}\mid g(\boldsymbol{x})<\tau\}.$ Then, we focus on the selective risk corresponds to the samples $\tilde{x}$ drawn from the distribution
$\tilde{\boldsymbol{x}}\sim P(\boldsymbol{x}\mid g(\boldsymbol{x})<\tau):=P_{\mathcal{X}_{\tau}},$ (8)
which leads to the selective area under risk converage curve ($\text{AU}_{P_{\mathcal{X}_{\tau}}}$):
$\text{AU}_{P_{\mathcal{X}_{\tau}}}(f)=\mathbb{E}_{\tilde{\boldsymbol{x}}\sim P_{\mathcal{X}_{\tau}}}\left[\frac{\mathbb{E}_{P(x,y)}\left[\ell(f(\boldsymbol{x}),\boldsymbol{y})\mathbb{I}[g(\boldsymbol{x})\geq g(\tilde{\boldsymbol{x}})]\right]}{\mathbb{E}_{\boldsymbol{x}^{\prime}\sim P_{x}}\mathbb{I}[g(\boldsymbol{x}^{\prime})\geq g(\tilde{\boldsymbol{x}})]}\right$]. (9)
Noticing that the expectation in the numerator can be swapped with the expectation outside, the equation above can then be written as:
$\text{AU}_{P_{\mathcal{X}_{\tau}}}(f)=\mathbb{E}_{\text{P}(\boldsymbol{x},\boldsymbol{y})}\left[\beta(\boldsymbol{x})\ell\left(f\left(\boldsymbol{x}\right),\boldsymbol{y}\right)\right]$ (10)
where
[…]
Proposition 4.6 (An Equivalent Expression for $\beta(x)$).
Given the CDF of the CSF $g$ with its inverse denoted as $g^{-1}$, the expression for $\beta(\boldsymbol{x})$ from Eq. (11) is equivalent to
$\beta(\boldsymbol{x})=\begin{cases}-\ln(1-G(g^{-1}(\tau)))&\text{if }\boldsymbol{x}\in\mathcal{X}\setminus\mathcal{X}_{\tau},\\ -\ln(1-G(\boldsymbol{x}))&\text{if }\boldsymbol{x}\in\mathcal{X}_{\tau}.\end{cases}$ (12)
Proof.
See Appendix B.2. ∎
[Bin-based CDF Approximation.] Let $G(\boldsymbol{x})$ denote the CDF of the CSF, as defined in (6), under the data distribution. We approximate $G$ using a binning-based approach. Given bin edges $\mathcal{E}=\{e_{0},e_{1},\ldots,e_{K}\}$ with $e_{0}<\cdots<e_{K}$, the corresponding bin centers are defined as:
$c_{k}=\frac{e_{k}+e_{k+1}}{2},\quad k=0,\dots,K-1.$ (13)
For a set of confidence scores $\{g(\boldsymbol{x}_{i})\}_{i=1}^{n}$, let $h_{k}$ denote the count in each bin:
$h_{k}=\sum_{i=1}^{n}\mathbb{I}[e_{k}\leq g(\boldsymbol{x}_{i})<e_{k+1}].$ (14)
To maintain differentiability for backpropagation, we approximate the CDF using a smoothed step function based on the sigmoid function. The estimated CDF at any point $\boldsymbol{x}$ is:
$\widehat{G}_{\nu}(\boldsymbol{x})=\frac{1}{n}\sum_{k=0}^{K-1}h_{k}\,\sigma\!\left(\frac{\boldsymbol{x}-c_{k}}{\nu}\right),\quad\sigma(t)=\frac{1}{1+e^{-t}},$ (15)
where $\nu>0$ is the smoothing parameter. A detailed discussion of this parameter is provided in Appendix E.2.
[Finite-Sample Estimator.] Given a threshold $\tau$ defining the low-confidence region $\mathcal{X}_{\tau}$, we estimate the sample weights in Eq. (12) as:
$\widehat{\beta}_{i}=\begin{cases}-\ln\!\big(1-\widehat{G}_{\nu}(g^{-1}(\tau))\big),&g(\boldsymbol{x}_{i})\geq\tau,\\[4.30554pt] -\ln\!\big(1-\widehat{G}_{\nu}(\boldsymbol{x}_{i})\big),&g(\boldsymbol{x}_{i})<\tau.\end{cases}$ (16)
This weighting assigns uniform weights to high-confidence samples and rank-based weights to low-confidence samples. Substituting (16) into Eq. (10) yields the differentiable finite-sample estimator:
$\widehat{\mathrm{AU}}_{P_{\mathcal{X}_{\tau}}}(f)=\frac{1}{n}\sum_{i=1}^{n}\widehat{\beta}_{i}\cdot\ell\big(f(\boldsymbol{x}_{i}),\boldsymbol{y}_{i}\big).$ (17)
Algorithm 1 Optimization with $\widehat{\mathrm{AU}}_{P_{\mathcal{X}_{\tau}}}$ loss
@@ 4.2 Selective Risk Loss
1:Input: training data $\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\}_{i=1}^{n}$, model $f_{\theta}$, CSF $g(\cdot)$, bin edges $\mathcal{E}$, smoothing parameter $\nu$, low-confidence threshold $\tau$, learning rate $\eta$
2:Forward Pass:
3:  $\boldsymbol{p}_{i}\leftarrow f_{\theta}(\boldsymbol{x}_{i})$
4:  Update $\widehat{G}_{\nu}(\boldsymbol{x})$ with $g(\boldsymbol{p}_{i})$ via Eq. (15)
5:  Compute $\widehat{\beta}_{i}$ via Eq. (16)
6:  $\widehat{\mathrm{AU}}\leftarrow\frac{1}{n}\sum_{i=1}^{n}\widehat{\beta}_{i}\cdot\ell(\boldsymbol{p}_{i},\boldsymbol{y}_{i})$ {Eq. (17)}
@@
9:  Compute $\nabla_{\boldsymbol{p}_{i}}\widehat{\mathrm{AU}}$ via Eq.(27)
10:  $\nabla_{\theta}\widehat{\mathrm{AU}}\mathrel{+}=(\nabla_{\boldsymbol{p}_{i}}\widehat{\mathrm{AU}})^{\top}\nabla_{\theta}\boldsymbol{p}_{i}$
@@
12:Parameter Update: $\theta\leftarrow\theta-\eta\nabla_{\theta}\widehat{\mathrm{AU}}$
Unlike the original AURC computation, which requires sorting all $n$ samples based on the CSF and thus incurs a complexity of $\mathcal{O}(n\log n)$, the proposed bin-based loss does not require any sorting. Instead, it approximates the CDF of confidence scores using $K$ bins, resulting in a computational cost of $\mathcal{O}(nK)$, which is typically lower than that of AURC and generally more efficient than kernel-based calibration error losses, which require quadratic time complexity. As with AURC, the behavior of this loss depends on the choice of CSF $g$, and, as we show in later sections, different CSFs may lead to different calibration performance.
(a) CIFAR-10
(b) CIFAR-100
(c) Tiny-ImageNet
Figure 2: Test ECE of $\widehat{\mathrm{AU}}_{P_{\mathcal{X}_{\tau}}}$ loss with different CSFs, reported as mean $\times 10^{2}$.
@@ 5 Experiments
Table 1: Test accuracy and ECE ($\downarrow$), reported as mean $\times 10^{2}$.
Model XE FL-53 Inv.FL Dual KDE-XE AURC $\mathbf{\text{AU}_{P_{\mathcal{X}_{\tau}}}}$ XE FL-53 Inv.FL Dual KDE-XE AURC $\mathbf{\text{AU}_{P_{\mathcal{X}_{\tau}}}}$
CIFAR-10    Accuracy ($\uparrow$) ECE ($\downarrow$)
@@
ResNet110 94.47 93.80 93.92 93.72 94.68 93.05 93.39 3.04 4.44 4.80 2.08 2.80 2.44 1.76
[…]
PreResNet56 93.97 92.97 94.14 93.72 94.09 92.63 92.79 2.62 5.50 3.89 2.07 2.57 1.75 0.88
CIFAR-100    Accuracy ($\uparrow$) ECE ($\downarrow$)
[…]
ResNet110 75.27 77.48 77.36 76.30 78.27 74.33 74.11 7.29 4.57 10.90 4.60 9.09 5.23 2.73
[…]
PreResNet56 74.28 72.93 74.18 72.96 74.69 71.91 73.89 7.10 2.44 9.81 0.96 7.42 4.88 1.81
Tiny-ImageNet Accuracy ($\uparrow$) ECE ($\downarrow$)
[…]
Swin-Tiny 85.36 84.31 85.64 85.42 85.36 82.35 84.38 1.62 7.69 3.00 1.60 1.71 2.02 1.45
Dataset Model XE FL-53 Inv. FL Dual KDE-XE AURC $\mathbf{\text{AU}_{P_{\mathcal{X}_{\tau}}}}$
[…]
Table 2: Mean cwECE $\times 10^{2}$ ($\downarrow$) for original models and after applying post-hoc calibration methods including TS. Temperature values are included in parentheses, and the temperature $T$ is selected by minimizing ECE on the training set.
0.00.20.40.60.81.00.00.20.40.60.81.0ConfidenceAccuracyPerfect CalibrationECE=5.83
(a) FL-53
0.00.20.40.60.81.00.00.20.40.60.81.0ConfidenceAccuracyPerfect CalibrationECE=3.92